OpenStudy (anonymous):

What is the solution set of |x + 3| - 2 = -5

mathslover (mathslover):

First add 2 both sides. |x+3| = -5+ 2 => |x+3| = -3

OpenStudy (anonymous):

a) -6 or 0 b) No solution c) -5 or 5 d) -3 or 3

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

so the answer is d. correct?

OpenStudy (anonymous):

oh i see what i did wrong.

mathslover (mathslover):

Oh wait @danny562 , its "no solution"

mathslover (mathslover):

Sorry didn't notice that. |anything| \(\ne\) (Negative) Absolute value of anything can not be negative.

mathslover (mathslover):

Hence it is B) NO Solution.

OpenStudy (anonymous):

oh oh well to late now its cool bro.